OpenStudy (itsmichelle29):

Using the graph of f(x) = log2x below, approximate the value of y in the equation 22y = 4.

OpenStudy (itsmichelle29):

2^2y=4

OpenStudy (michele_laino):

if we take the logarithm with base 2 of your equation, we get: \[\Large \begin{gathered} {2^{2y}} = 4 \hfill \\ \hfill \\ 2y{\log _2}2 = {\log _2}4 \hfill \\ \hfill \\ 2y = {\log _2}4 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

now, using the properties of logarithm in any base, we can write: \[\Large {\log _2}4 = {\log _2}\left( {{2^2}} \right) = 2{\log _2}2\]

OpenStudy (itsmichelle29):

so the answer would be 1

OpenStudy (michele_laino):

yes!

OpenStudy (itsmichelle29):

thank you so much

OpenStudy (michele_laino):

:)